# AdvDiff: Generating Unrestricted Adversarial Examples using Diffusion Models

### v3

@@ Abstract
Unrestricted adversarial attacks present a serious threat to deep learning models and adversarial defense techniques. They pose severe security problems for deep learning applications because they can effectively bypass defense mechanisms. However, previous attack methods often utilize Generative Adversarial Networks (GANs), which are not theoretically provable and thus generate unrealistic examples by incorporating adversarial objectives, especially for large-scale datasets like ImageNet. In this paper, we propose a new method, called AdvDiff, to generate unrestricted adversarial examples with diffusion models. We design two novel adversarial guidance techniques to conduct adversarial sampling in the reverse generation process of diffusion models. These two techniques are effective and stable to generate high-quality, realistic adversarial examples by integrating gradients of the target classifier interpretably. Experimental results on MNIST and ImageNet datasets demonstrate that AdvDiff is effective to generate unrestricted adversarial examples, which outperforms GAN-based methods in terms of attack performance and generation quality.
@@ 1 Introduction
While the deep learning (DL) community continues to explore the wide range of applications of DL models, researchers [37] have demonstrated that these models are highly susceptible to deception by adversarial examples. Adversarial examples are generated by adding perturbations to clean data. The perturbed examples can deceive DL classifiers with high confidence while remaining imperceptible to humans. Many strong attack methods [24, 12, 3, 7] are proposed and investigated to improve the robustness of DL models.
@@
Diffusion models [16] are likelihood-based generative models proposed recently, which emerged as a strong competitor to GANs. Diffusion models have outperformed GANs for image synthesis tasks [11, 31, 19]. Compared with GAN models, diffusion models are more stable during training and provide better distribution coverage. Diffusion models contain two processes: a forward diffusion process and a reverse generation process. The forward diffusion process gradually adds Gaussian noise to the data and eventually transforms it into noise. The reverse generation process aims to recover the data from the noise by a denoising-like technique. A well-trained diffusion model is capable of generating images with random noise input. Similar to GAN models, diffusion models can achieve conditional image generation with given label information.
@@
GAN-based unrestricted adversarial attacks often exhibit poor performance on high-quality datasets, particularly in terms of visual quality, because of GAN’s weak interpretability and unstable training. These attacks tend to generate low-quality adversarial examples compared to benign GAN examples. Therefore, these attacks are not imperceptible among GAN synthetic data. Diffusion models, however, offer state-of-the-art generation performance [11] on challenging datasets like LSUN [43] and ImageNet [9]. The conditional diffusion models can generate images based on specific conditions by sampling from a perturbed conditional Gaussian noise, which can be carefully modified with adversarial objectives. These properties make diffusion models more suitable for conducting unrestricted adversarial attacks.
@@
In this paper, we propose a novel unrestricted adversarial attack method called AdvDiff that utilizes diffusion models for adversarial examples generation, as shown in Figure 1. Specifically, AdvDiff uses a trained conditional diffusion model to conduct adversarial attacks with two new adversarial guidance techniques. 1) During the reverse generation process, we gradually add adversarial guidance by increasing the likelihood of the target attack label. 2) We perform the reverse generation process multiple times, adding adversarial prior knowledge to the initial noise with the noise sampling guidance. Our theoretical analysis indicates that these adversarial guidance techniques can effectively craft adversarial examples by the reverse generation process with adversarial conditional sampling. Furthermore, the sampling of AdvDiff benefits from stable and high sample quality of the diffusion models sampling, which leads to the generation of realistic unrestricted adversarial examples. Through extensive experiments conducted on two datasets, i.e., the high-quality dataset ImageNet, and the small, robust dataset MNIST, we have observed a significant improvement in the attack performance using AdvDiff with diffusion models. These results prove that our proposed AdvDiff is more effective than previous GAN-based methods in conducting unrestricted adversarial attacks to generate high-fidelity and diverse examples without decreasing the generation quality.
@@
Figure 1: The two new guidance techniques in our AdvDiff to generate unrestricted adversarial examples. During the reverse generation process, the adversarial guidance is added at timestep $x^{t}$, which injects the adversarial objective $y_{a}$ into the diffusion process. The noise sampling guidance modifies the original noise by increasing the conditional likelihood of $y_{a}$.
@@ 2.1 Diffusion Model
The Denoising Diffusion Probabilistic Model (DDPM) [16] is utilized to learn the Gaussian transitions from $p(x_{T})=\mathcal{N}(x_{T};0,\mathbf{I})$ to recover the data $x_{0}\sim q(x_{0})$ with a Markov chain. We call this denoising-like process the reverse generation process. Following the pre-defined $T$ time steps, DDPM obtains a sequence of noisy data $\{x_{1},\dots,x_{N-1}\}$ and finally recovers the data $x_{0}$. It is defined as:
@@
Training of the DDPM requires accurate inference of the mean value $\mu_{\theta}$ with a deep learning model in the reverse generation process. The objective is to learn the variational lower-bound (VLB) on $\log p_{\theta}(x_{0})$. In order to complete the training objective, we train the model $\varepsilon_{\theta}$ to predict the added Gaussian noise in the forward diffusion process. The standard mean-squared error loss is adopted:
@@ 2.2 Classifier-Guided Guidance
$\hat{\mu}_{\theta}(x_{t},t)=\mu_{\theta}(x_{t},t)+s\cdot\Sigma_{\theta}(x_{t}|y)\nabla_{x_{t}}\log p_{\phi}(y|x_{t})$ (4)
@@ 3.1 Rethinking Unrestricted Adversarial Examples
However, due to GAN’s poor interpretability, there’s no theoretical support on $z_{adv}$ that can craft UAEs with normally trained GANs. The generator of GAN is not trained with $z+\epsilon\nabla\mathcal{L}$ but only $z\sim\mathcal{N}(0,\textbf{I})$. Therefore, GAN-based UAEs encounter a significant decrease in generation quality because samples with $z_{adv}$ are not well-trained compared with samples with $z\sim\mathcal{N}(0,\textbf{I})$. Moreover, the GAN latents are sampled from low dimensional latent spaces. Therefore, GANs are extremely sensitive to the latent $z$ [34, 22]. If we inject gradients of the classification results into GAN latents, GAN-based methods are more likely to generate flipped-label UAEs (images corresponding to the targeted attack label $y_{a}$ instead of the conditional generation label $y$) and distorted UAEs. However, these generation issues are hard to address only by attack success rate (ASR). In other words, even with a high ASR, some of the successful UAEs with GAN-based methods should be identified as failure cases for poor visual quality. However, such cases can not be reflected by ASR but can be evaluated by generation quality. All these problems may indicate that GAN models are not suitable for generative-based adversarial attacks.
@@
3.2 General Adversarial Diffusion Sampling
@@ 3.5 Training-Free Adversarial Attack
The proposed adversarial attack does not require additional modification on the training of the diffusion model. The adversarial examples are sampled by using Algorithm 1 over the trained classifier-free diffusion model $\epsilon_{\theta}(\cdot)$.
@@
The adversarial diffusion sampling starts with a random noisy input $x_{T}$. After the classifier-free guidance sampling [15] at step 6, we utilize the adversarial guidance to incorporate the adversarial gradient into the sampling process. After the sampling, we can obtain the $x_{0}$ with the adversarial guidance. To achieve more success and stable adversarial sampling, we can integrate the gradient of the target model’s classification of $x_{0}$ to the beginning noisy input $x_{T}$ for $N$ iterations. Finally, we get the adversarial examples that fool the target model with the label $y_{a}$. We give the AdvDiff algorithm on DDIM in the Appendix.
4 Related Work
Since Szegedy et al. [37] had proved that DL models are extremely vulnerable to adversarial attacks, researchers have been digging into improving the model’s adversarial robustness by proposing stronger adversarial attack methods and their counter-measurements.
Perturbation-based adversarial examples with generative models: Most related works performed adversarial attacks by perturbing a subset of clean data to fool the target classifier. These attacks [3, 24, 20] attempted to generate better perturbations with higher attack success rates and smaller perturbations. With the emergence of generative models, end-to-end adversarial attacks [1, 41, 28] have greatly improved the generation efficiency by pre-training the generative module. These methods integrate the advertorial loss into the training of generative models and generate adversarial examples by trained generators with clean data.
Unrestricted adversarial examples with generative models: Perturbation-based adversarial examples require insignificant norm distance to the given clean data in order to guarantee the indistinguishability, which only covers a small fraction of all possible adversarial examples [36]. To remove such restrictions, Song et al. [36] proposed an unrestricted adversarial attack method that searches over the latent space of the input noise vector with an adversarial loss function and a well-trained AC-GAN [27]. Inspired by Song’s work [36], recent works [29, 40] made improvements over the generation quality and generation efficiency of UAEs. However, the performance of existing approaches suffers from the unstable training of GAN models as well as the lack of theoretical support of GAN-based UAEs. Therefore, we provide an effective and theoretically analyzed solution with the diffusion model in this paper.
@@ 5 Experiments
Comparisons. It is not applicable to give a clear comparison between perturbation-based attacks and unrestricted attacks because perturbation-based attacks have the corresponding ground truth while unrestricted attacks do not. We mainly compare our method with the GAN-based unrestricted adversarial attack U-GAN [36] and give the discussion with the PGD [24], BIM [12], and C&W [3] perturbation-based attacks. For U-GAN, We adopt AC-GAN [27] for the MNIST dataset, and SAGAN [45] and BigGAN [2] for the ImageNet dataset, as AC-GAN has shown poor performance on ImageNet. Additionally, we use PGD gradient to replace the adversarial guidance in our method for a clear comparison, which is represented by “AdvDiff-PGD”.
Implementation details. Because our adversarial diffusion sampling does not require additional training to the original diffusion model, we use the pre-trained diffusion model in our experiment. We adopt DDPM [16] with classifier-free guidance for the MNIST dataset and Latent Diffusion Model (LDM) [31] with DDIM sampler for the ImageNet dataset. For MNIST dataset, we use $N=10$, $s=0.5$, and $a=1.0$, And $N=5$, $s=0.25$, and $a=0.5$ for ImageNet dataset. More details and experiments are given in the Appendix.
Evaluation metrics. We utilize the top-1 classification result to evaluate the Attack Success Rate (ASR) on different attack methods. As discussed earlier, GAN-based UAEs often encounter severe generation quality drops compared to benign GAN samples. Therefore, we give comparisons of generation performance on ImageNet to evaluate the attack performance of different UAEs in imperceptibly. The results are averaged with five runs. We use ResNet50 as the target model for default settings.
Figure 4: Comparisons of unrestricted adversarial attacks between GANs and diffusion models on two datasets.. Left: generated samples from U-GAN (BigGAN for ImageNet dataset). Right: generated samples from AdvDiff. We generate unrestricted adversarial examples on the MNIST “0” label and ImageNet “mushroom” label. U-GAN is more likely to generate adversarial examples with the target label, i.e., examples with red font. However, AdvDiff tends to generate the “false negative” samples by the target classifier by combing features from the target label.
@@ 5.1 Attack Performance
Table 1: The attack success rate on MNIST dataset. The unrestricted adversarial examples effortlessly break through current state-of-the-art defense methods, the reason is that unrestricted adversarial attacks adopt a different threat model.
@@
ImageNet It is reported that deep learning models on ImageNet are extremely vulnerable to adversarial attacks. However, the state-of-the-art adversarial defense DiffPure [26] and adversarial training [24] can still defend against around 50% of the perturbation-based attacks, as reported in Table 2. Most of UAEs evade the current defenses, but the generation quality of U-GAN is relatively poor compared to our adversarial examples. This phenomenon also shows that the performance of UAEs is heavily affected by the generation quality of the generation model. The adversarial examples generated by AdvDiff are more aggressive and stealthy than U-GAN’s.
@@
PGD 99.8 51.2 56.9 99.5 45.3 50.6
U-SAGAN 99.3 95.9 98.5 98.9 90.4 97.6
U-BigGAN 96.8 89.2 94.3 96.5 84.5 91.2
AdvDiff-PGD 99.8 98.0 99.2 99.9 98.5 99.0
AdvDiff 99.8 98.1 99.3 99.9 98.5 99.1
@@ 5.2 Generation Quality: True ASR for UAEs
We witness similar ASR with U-GAN and AdvDiff. However, imperceptibility is also critical for a successful unrestricted adversarial attack, so we adopt the evaluation metrics in [11] to compare the generation quality with and without performing unrestricted attacks. Table 3 shows that the AdvDiff achieves an overwhelming better IS score and similar FID score on the large-scale ImageNet dataset, where FID [14] and IS [32] scores are commonly adopted for evaluating the quality of a generative model.

### v4

@@ Abstract
Unrestricted adversarial attacks present a serious threat to deep learning models and adversarial defense techniques. They pose severe security problems for deep learning applications because they can effectively bypass defense mechanisms. However, previous attack methods often directly inject Projected Gradient Descent (PGD) gradients into the sampling of generative models, which are not theoretically provable and thus generate unrealistic examples by incorporating adversarial objectives, especially for GAN-based methods on large-scale datasets like ImageNet. In this paper, we propose a new method, called AdvDiff, to generate unrestricted adversarial examples with diffusion models. We design two novel adversarial guidance techniques to conduct adversarial sampling in the reverse generation process of diffusion models. These two techniques are effective and stable in generating high-quality, realistic adversarial examples by integrating gradients of the target classifier interpretably. Experimental results on MNIST and ImageNet datasets demonstrate that AdvDiff is effective in generating unrestricted adversarial examples, which outperforms state-of-the-art unrestricted adversarial attack methods in terms of attack performance and generation quality.
Keywords:
Unrestricted Adversarial Attacks Diffusion Models Interpretable Adversarial Diffusion Sampling
@@ 1 Introduction
While the deep learning (DL) community continues to explore the wide range of applications of DL models, researchers [37] have demonstrated that these models are highly susceptible to deception by adversarial examples. Adversarial examples are generated by adding perturbations to clean data. The perturbed examples can deceive DL classifiers with high confidence while remaining imperceptible to humans. Many strong attack methods [25, 13, 3, 8, 22, 21] are proposed and investigated to improve the robustness of DL models.
@@
Diffusion models [17] are likelihood-based generative models proposed recently, which emerged as a strong competitor to GANs. Diffusion models have outperformed GANs for image synthesis tasks [12, 31, 19]. Compared with GAN models, diffusion models are more stable during training and provide better distribution coverage. Diffusion models contain two processes: a forward diffusion process and a reverse generation process. The forward diffusion process gradually adds Gaussian noise to the data and eventually transforms it into noise. The reverse generation process aims to recover the data from the noise by a denoising-like technique. A well-trained diffusion model is capable of generating images with random noise input. Similar to GAN models, diffusion models can achieve adversarial attacks by incorporating adversarial objectives [5, 6, 4].
@@
GAN-based unrestricted adversarial attacks often exhibit poor performance on high-quality datasets, particularly in terms of visual quality, because they directly add the PGD perturbations to the GAN latents without theoretic supports. These attacks tend to generate low-quality adversarial examples compared to benign GAN examples [35]. Therefore, these attacks are not imperceptible among GAN synthetic data. Diffusion models, however, offer state-of-the-art generation performance [12] on challenging datasets like LSUN [41] and ImageNet [10]. The conditional diffusion models can generate images based on specific conditions by sampling from a perturbed conditional Gaussian noise, which can be carefully modified with adversarial objectives. These properties make diffusion models more suitable for conducting unrestricted adversarial attacks. Nevertheless, existing adversarial attack methods using diffusion models [5, 6, 4] adopt similar PGD perturbations to the sample in each reverse generation process, making them generate relatively low-quality adversarial examples.
@@
In this paper, we propose a novel and interpretable unrestricted adversarial attack method called AdvDiff that utilizes diffusion models for adversarial examples generation, as shown in Figure 1. Specifically, AdvDiff uses a trained conditional diffusion model to conduct adversarial attacks with two new adversarial guidance techniques. 1) During the reverse generation process, we gradually add adversarial guidance by increasing the likelihood of the target attack label. 2) We perform the reverse generation process multiple times, adding adversarial prior knowledge to the initial noise with the noise sampling guidance.
Our theoretical analysis indicates that these adversarial guidance techniques can effectively craft adversarial examples by the reverse generation process with adversarial conditional sampling. Furthermore, the sampling of AdvDiff benefits from stable and high sample quality of the diffusion models sampling, which leads to the generation of realistic unrestricted adversarial examples. Through extensive experiments conducted on two datasets, i.e., the high-quality dataset ImageNet, and the small, robust dataset MNIST, we have observed a significant improvement in the attack performance using AdvDiff with diffusion models. These results prove that our proposed AdvDiff is more effective than previous unrestricted adversarial attack methods in conducting unrestricted adversarial attacks to generate high-fidelity and diverse examples without decreasing the generation quality.
@@
Figure 1: The two new guidance techniques in our AdvDiff to generate unrestricted adversarial examples. During the reverse generation process, the adversarial guidance is added at timestep $x_{t}$, which injects the adversarial objective $y_{a}$ into the diffusion process. The noise sampling guidance modifies the original noise by increasing the conditional likelihood of $y_{a}$.
@@ 2.1 Diffusion Model
The Denoising Diffusion Probabilistic Model (DDPM) [17] is utilized to learn the Gaussian transitions from $p(x_{t})=\mathcal{N}(x_{t};0,\mathbf{I})$ to recover the data $x_{0}\sim q(x_{0})$ with a Markov chain. We call this denoising-like process the reverse generation process. Following the pre-defined $T$ time steps, DDPM obtains a sequence of noisy data $\{x_{T-1},\dots,x_{1}\}$ and finally recovers the data $x_{0}$. It is defined as:
@@
Training of the DDPM requires accurate inference of the mean value $\mu_{\theta}$ with a deep learning model in the reverse generation process. The objective is to learn the variational lower-bound (VLB) on $\log p_{\theta}(x_{0})$. In order to complete the training objective, we train the model $\epsilon_{\theta}$ to predict the added Gaussian noise in the forward diffusion process. The standard mean-squared error loss is adopted:
@@ 2.2 Classifier-Guided Guidance
$\hat{\mu}_{\theta}(x_{t},t)=\mu_{\theta}(x_{t},t)+s\cdot\nabla_{x_{t}}\log p_{\phi}(y|x_{t})$ (4)
@@ 3.1 Rethinking Unrestricted Adversarial Examples
However, due to GAN’s poor interpretability, there’s no theoretical support on $z_{adv}$ that can craft UAEs with normally trained GANs. The generator of GAN is not trained with $z_{adv}=z+\nabla\mathcal{L}$ but only $z\sim\mathcal{N}(0,\textbf{I})$. Therefore, GAN-based UAEs encounter a significant decrease in generation quality because samples with $z_{adv}$ are not well-trained compared with samples with $z\sim\mathcal{N}(0,\textbf{I})$. Moreover, the GAN latents are sampled from low dimensional latent spaces. Therefore, GANs are extremely sensitive to the latent $z$ [33, 23]. If we inject gradients of the classification results into GAN latents, GAN-based methods are more likely to generate flipped-label UAEs (images corresponding to the targeted attack label $y_{a}$ instead of the conditional generation label $y$) and distorted UAEs. However, these generation issues are hard to address only by attack success rate (ASR). In other words, even with a high ASR, some of the successful UAEs with GAN-based methods should be identified as failure cases for poor visual quality. However, such cases can not be reflected by ASR but can be evaluated by generation quality. All these problems may indicate that GAN models are not suitable for generative-based adversarial attacks.
@@
3.2 Adversarial Diffusion Sampling with Theoretical Support
@@ 3.5 Training-Free Adversarial Attack
The proposed adversarial attack does not require additional modification on the training of the diffusion model. The adversarial examples are sampled by using Algorithm 1 over the trained classifier-free diffusion model $\epsilon_{\theta}(\cdot)$. We give the AdvDiff algorithm on DDIM in the Appendix.
@@ 4 Experiments
Comparisons. It is not applicable to give a clear comparison between perturbation attacks and unrestricted attacks because perturbation attacks have the corresponding ground truth while unrestricted attacks do not. We mainly compare our method with the unrestricted adversarial attack U-GAN [35] and give the discussion with the AutoAttack [8], PGD [25], BIM [13], and C&W [3] perturbation-based attacks under norm $\ell_{\text{inf}}=8/255$. For U-GAN, We adopt AC-GAN [29] for the MNIST dataset, and SAGAN [43] and BigGAN [2] for the ImageNet dataset, as AC-GAN has shown poor performance on ImageNet. We use the official code from DiffAttack [4] and implement AdvDiffuser by ourselves [5] for comparisons. We do not compare with Chen et al. [6], because they use a similar method as DiffAttack and without official code. Because existing diffusion model attacks are all untargeted attacks, we include the untargeted version of AdvDiff for a clear comparison, which is represented by “AdvDiff-Untargeted”.
Implementation details. Because our adversarial diffusion sampling does not require additional training to the original diffusion model, we use the pre-trained diffusion model in our experiment. We adopt DDPM [17] with classifier-free guidance for the MNIST dataset and Latent Diffusion Model (LDM) [31] with DDIM sampler for the ImageNet dataset. For MNIST dataset, we use $N=10$, $s=0.5$, and $a=1.0$, And $N=5$, $s=0.7$, and $a=0.5$ for ImageNet dataset. More details and experiments are given in the Appendix.
Evaluation metrics. We utilize the top-1 classification result to evaluate the Attack Success Rate (ASR) on different attack methods under untargeted attack settings. As discussed earlier, GAN-based UAEs often encounter severe generation quality drops compared to benign GAN samples. Therefore, we give comparisons of generation performance on ImageNet to evaluate the attack performance of different UAEs in imperceptibly. The results are averaged with five runs. We use ResNet50 as the target model for default settings.
Figure 4: Comparisons of unrestricted adversarial attacks between GANs and diffusion models on two datasets.. Left: generated samples from U-GAN (BigGAN for ImageNet dataset). Right: generated samples from AdvDiff. We generate unrestricted adversarial examples on the MNIST “0" label and ImageNet “mushroom" label. U-GAN is more likely to generate adversarial examples with the target label, i.e., examples with red font. However, AdvDiff tends to generate the “false negative" samples by the target classifier by combing features from the target label.
@@ 4.1 Attack Performance
Table 1: The attack success rate on MNIST dataset.
@@
AutoAttack 95.1 22.2 56.2 94.9 20.6 55.4 0.5
U-SAGAN 99.3 30.5 80.6 98.9 28.6 70.1 10.4
U-BigGAN 96.8 40.1 81.5 96.5 35.5 78.4 11.2
AdvDiffuser 95.4 28.9 90.6 94.6 26.5 88.9 38.6
DiffAttack 92.8 30.6 88.4 90.6 27.6 85.3 28.2
AdvDiff 99.8 41.6 92.4 99.9 38.5 90.6 9.2
AdvDiff-Untargeted 99.5 75.2 94.5 99.4 70.5 92.6 9.6
ImageNet It is reported that deep learning models on ImageNet are extremely vulnerable to adversarial attacks. However, the state-of-the-art adversarial defense DiffPure [28] and adversarial training [25] can still defend against the perturbation-based attacks, as reported in Table 2. More UAEs evade the current defenses, but the generation quality of U-GAN is relatively poor compared to our adversarial examples. This phenomenon also shows that the performance of UAEs is heavily affected by the generation quality of the generation model. The adversarial examples generated by AdvDiff are more aggressive and stealthy than U-GAN’s. Meanwhile, the generation speed of AdvDiff is the best among all the unrestricted adversarial attack methods. Note that we adopt the clean images generated by LDM to achieve DiffAttack and AutoAttack for a fair comparison.
@@ 4.2 Generation Quality: True ASR for UAEs
We witness similar ASR with U-GAN and AdvDiff. However, imperceptibility is also critical for a successful unrestricted adversarial attack, so we adopt the evaluation metrics in [12] to compare the generation quality with and without performing unrestricted attacks. Table 3 shows that the AdvDiff achieves an overwhelming better IS score and similar FID score on the large-scale ImageNet dataset, where FID [16] and IS [32] scores are commonly adopted for evaluating the quality of a generative model. Because the generation of UAEs does not modify the data distribution of the generated images, the Precision score can be inferred as generation quality, while the Recall score indicates the flipped-label problems. We witness the frequent generation of flipped-label UAEs and low-quality UAEs from GAN-based methods, which is reflected by the decrease in the Precision score and the increase in the Recall score. Figure 4 illustrates this problem with some examples. It can be further proved that U-BigGAN achieves much higher image quality on non-reference metrics than reference metrics, as shown in Table 4.
@@
AdvDiff-Untargeted 22.8 0.14 0.85 16.2 76.8
We find the IS score is heavily affected by the transferability of adversarial examples due to the calculation method. Therefore, we further compare the image quality of adversarial examples by commonly used metrics in Table 4. The results show that AdvDiff (average 5 out of 5) and AdvDiff-Untargeted (average 4 out of 5) outperform existing adversarial attack methods using diffusion models. The perturbation-based adversarial attacks, i.e., AutoAttack, achieve much worse image quality compared with UAEs.
4.3 UAEs against Defenses and Black-box Models
Current defenses assume the adversarial examples are based on perturbations over data from the training dataset, i.e., $x_{adv}=x+\nabla\mathcal{L},x\in D$. However, UAEs are synthetic data generated by the generative model. Because of different data sources, current defenses are hard to defend UAEs, which brings severe security concerns to deep learning applications. The proposed AdvDiff achieves an average of 36.8% ASR against various defenses, while AutoAttack only achieves 30.7% ASR with significantly lower image quality. We also test the attack transferability of AdvDiff and the results show that the untargeted version of AdvDiff achieves the best performance against black-box models. Experiment results are given in Table 5.
  Method ResNet-152 [15] Inception v3 [36] ViT-B [14] BEiT [1]
AutoAttack 32.5 38.6 9.3 45.3
U-BigGAN 30.8 35.3 30.1 69.4
AdvDiffuser 18.3 20.0 18.5 79.4
DiffAttack 21.1 43.9 17.4 78.0
AdvDiff 20.5 14.9 17.8 78.8
AdvDiff-Untargeted 52.0 42.7 36.0 81.5
  Method Adv-Inception [25] AdvProp [38] DiffPure [28] HGD [24]
AutoAttack 14.6 69.6 22.2 20.5
U-BigGAN 40.6 75.2 40.1 22.6
AdvDiffuser 24.4 84.0 30.5 10.8
DiffAttack 30.9 85.1 30.6 20.5
AdvDiff 19.4 89.7 41.6 17.8
AdvDiff-Untargeted 60.1 95.3 75.2 53.8
  Method R&P [39] RS [7] NRP [27] Bit-Red [40]
AutoAttack 20.6 38.9 39.4 19.8
U-BigGAN 14.2 34.5 30.9 13.1
AdvDiffuser 15.4 38.4 40.5 11.4
DiffAttack 23.7 40.8 38.5 20.1
AdvDiff 17.4 47.6 45.2 15.8
AdvDiff-Untargeted 56.8 82.8 74.2 52.6
4.4 Better Adversarial Diffusion Sampling
We present detailed comparisons with DiffAttack and AdvDiffuser. The results show that the proposed adversarial guidance achieves significantly higher generation quality than PGD-based adversarial guidance. With PGD gradient guidance, the diffusion model generates images with a similar Recall score but a much lower Precision score, which indicates that the PGD gradient influences the benign generation process and causes the generation of low-quality images. The result proves that the adversarial guidance of diffusion models should be carefully designed without affecting the benign sampling process. Meanwhile, the generation speed of AdvDiff is the best among the existing diffusion attack methods. Note that AdvDiff (36.8%) sightly outperforms AdvDiffuser (32.0%) and DiffAttack (36.2%) against defenses. However, previous attacks achieve slightly better transfer attack performance than the original AdvDiff. The reason could be the gradient of the cross-entropy loss is shared among nearly all the deep learning models and is better at attack transferability against these models. Nevertheless, the untargeted version of AdvDiff achieves overwhelmingly better performance, which further demonstrates the effectiveness of the proposed adversarial sampling. But the generation quality is affected, we leave a better design in the future work.